# Nearly Zero-Cost Protection Against Mimicry by Personalized Diffusion Models

### v1

@@ 1 Introduction
Large-scale diffusion models [song2020score, ho2020denoising, rombach2022high, podell2023sdxl] have achieved remarkable success in the realm of image synthesis task, revolutionizing the way we create and manipulate digital imagery [zhang2023adding, kim2023diffblender, huang2023composer, meng2021sdedit, mou2023t2i]. By adopting recently emerged personalization techniques, it is now possible to develop diffusion models that can generate images in one’s unique styles [ruiz2023dreambooth, gal2022image, kumari2023multi, sohn2023styledrop, ahn2023dreamstyler]. However, the very capabilities that make these models so valuable also pose significant risks. Diffusion models can allow malicious users to replicate an individual’s artwork without consent, easily stealing their creative output [shan2023glaze]. Furthermore, diffusion models are exceedingly adept at mimicking faces and objects [ruiz2023dreambooth, kumari2023multi]. This proficiency is not limited to benign applications but can be extended to the making of deepfakes or fake news. The implications for societal harm are profound, as these technologies could be used to undermine public trust, distort political discourse, and violate personal privacy.
@@
Current efforts to protect images against diffusion models have evolved through adversarial perturbation [goodfellow2014explaining]. By introducing perturbations to a given image, they render the protected image resistant to diffusion-based personalization methods, causing the diffusion model to generate distorted outputs. Existing methods primarily rely on two approaches: minimizing the distance between the latent features of the protected and a pre-defined target image within a VAE encoder (e.g. PhotoGuard [salman2023raising] and Glaze [shan2023glaze]), or fooling the noise prediction network (e.g. AdvDM [liang2023adversarial] and Anti-DB [van2023anti]). Despite numerous advancements, all current frameworks fundamentally start from an approach that optimizes perturbations during inference [madry2017towards] when protection is requested, casting doubt on the practicality of these methods to be used in many real applications.
@@
For a protection framework to become a practical solution, the following conditions should be satisfied: (1) Protection Efficacy: Obviously, image protection method should effectively counteract against diffusion-based personalization. Many recent studies have aimed to maximize this aspect by analyzing the internal dynamics of diffusion models [xue2023toward, zheng2023understanding]. (2) Invisibility: Although some studies strive to be as invisible as possible, adversarial perturbations inevitably leave visible traces on images. However, for commercial services, minimizing these traces is crucial. Moreover, as analyzed by ahn2024imperceptible, the task of preventing mimicry tends to leave more noticeable traces, particularly in images with flat textures (e.g. cartoon or illustration). Despite its importance, only a handful of studies highlight this issue [ahn2024imperceptible]. (3) Latency: All existing methods update perturbations during the inference phase through iterative optimization. However, this process is inherently time-consuming; e.g. protecting a $512^{2}$-px image can take up to 5-120 minutes on a CPU and 7-200 seconds even on a high-end A100 GPU. The extensive time required by current protection methods poses a significant barrier that prevents ordinary users from utilizing image protection, thereby still exposing them to the risk of misuse of generative models. Consequently, in order to democratize this technology, it is crucial to develop protection framework that can operate effectively on less powerful devices. Unfortunately, despite extensive progresses in the image protection task, there is scant focus on reducing latency.
@@
In this work, we propose a novel protection framework, FastProtect, which meets all the requirements, with a particular focus on latency. FastProtect diverges from prevalent approaches by adopting a pre-trained perturbation. Although universal adversarial perturbation (UAP) [moosavi2017universal] is proposed in adversarial attack, we observed that it significantly reduces protection efficacy for our task. To counteract this, we introduce a mixture-of-perturbations (MoP). Unlike UAP, which relies on a single perturbation, MoP prepares multiple perturbations and selects one of them based on the input image’s latent code. We also propose a multi-layer protection (MLP) loss that utilizes the intermediate features when calculating the protection loss. It intensifies the distortion when the protected images go through fine-tuning by personalization methods, thus enhancing the protection efficacy without any additional cost at inference.
@@
For the inference, we introduce an adaptive targeted protection to enhance protection efficacy with minimal computational overhead. Our observations suggest that in targeted protection, which minimizes the distance between the input image’s latent code and that of a (pre-defined) target, selecting an appropriate target image significantly influences protection performance. Taking this into account, we adaptively determine the best target image based on the input image’s latent code and then apply the perturbation trained for that specific target. To enhance invisibility, we propose an adaptive protection strength approach using the LPIPS [zhao2023unlearnable] distance. Unlike ahn2024imperceptible that leverages multiple just-noticeable difference (JND) maps, our approach does not require ad-hoc modules nor does it significantly increase computational load; it only involves an additional forward pass of LPIPS (i.e. AlexNet [krizhevsky2012imagenet]).
@@ 2 Background
Diffusion Models. These have gained prominence due to their ability to generate high-quality images. Latent Diffusion Model (LDM) [rombach2022high] is particularly highligted for its exceptional quality and efficiency. Within LDM, a VAE encoder $\mathcal{E}$, transforms an input image $\mathbf{x}$ into a latent code $\mathbf{z}=\mathcal{E}(\mathbf{x})$. This code is then reconstructed back into the image domain by a decoder $\mathcal{D}$; $\mathbf{x}^{\prime}=\mathcal{D}(\mathbf{z}^{\prime})$. The diffusion process makes a latent code $\mathbf{z}^{\prime}$ by incorporating external factors, $y$, such as textual prompts. The training of LDM is driven by a denoising loss function at each timestep $t$ as:
@@
where a denoising network $\epsilon_{\theta}$ restores the noised latent code $\mathbf{z}_{t}$ based on the timestep $t$ and a conditioning vector $c(y)$. Recent advancements explore the personalization (or few-shot fine-tuning) of LDM with a few reference images through two primary approaches: textual inversion [gal2022image], which utilizes the embedding space of CLIP [radford2021learning] while keeping the denoising network fixed, and model optimization [ruiz2023dreambooth], which directly updates the denoising network.
Protection Against Diffusion-based Mimicry. Current protection frameworks apply adversarial perturbations $\delta$ to image $\mathbf{x}$, producing a protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$ using projected gradient descent (PGD) [madry2017towards]. With a protection loss $\mathcal{L_{P}}$, current methods obtain the protected image of $i$-th optimization step by a signed gradient ascent with step function sgn and step length $\alpha$ as given by
@@
For the objective function, semantic or texture losses are mostly used [liang2023mist]. The semantic loss, $\mathcal{L_{S}}(\mathbf{x})=\mathcal{L_{SD}}(\mathbf{x})$, is designed to disrupt the denoising process of LDM, misleading it to generate samples that deviate from the original images. AdvDM [liang2023adversarial] adopts this loss and Anti-DB [van2023anti] is improved to be more robust against DreamBooth [ruiz2023dreambooth] by incorporating its update process within the optimization. The texture loss, $\mathcal{L_{T}}(\mathbf{x})=-||\mathcal{E}(\mathbf{x})-\mathcal{E}(\mathbf{y})||^{2}_{2}$, aims to pushing the latent code of $\mathbf{x}$ towards the target image’s latent code. Glaze [shan2023glaze] and PhotoGuard [salman2023raising] belong to this category. Upon these objective functions, We can now define a universal loss as $\mathcal{L_{P}}=\lambda_{\mathcal{S}}\mathcal{L}_{S}+\lambda_{\mathcal{T}}\mathcal{L}_{T}$, where $\{\lambda_{\mathcal{S}},\lambda_{\mathcal{T}}\}\geq 0$ are balancing factors. Mist [liang2023mist] utilizes this universal loss function to effectively capture the characteristics of both objectives. Diff-Protect [xue2023toward], building on Mist, introduces a score distillation trick to achieve more efficient protection. Impasto [ahn2024imperceptible] incorporates perceptual-oriented components designed for imperceptible protection.
@@
Since all the current protection frameworks leverage iterative optimization during inference, they can yield image-specific and high-performing perturbations [salman2023raising, liang2023mist, shan2023glaze, liang2023adversarial, ahn2024imperceptible, van2023anti, xue2023toward, zheng2023understanding]. Despite this advantage, a significant drawback is the substantial time required to operate protection. For instance, to protect a $512^{2}$-px image, even the fastest model (e.g. PhotoGuard [salman2023raising]) requires around 7 seconds on an A100 GPU and 350 seconds on a CPU. Consequently, we explore alternatives to this iterative process to enhance efficiency without compromising the protection performance.
@@
Figure 2: Model overview. (a) Current iterative optimization approaches lack a training phase and perform optimization during inference, resulting in extremely slow protection. (b) UAP [moosavi2017universal] introduces pre-training of perturbations, but their image-agnostic nature leads to degraded protection efficacy. (c) Combining the advantages of both paradigms, FastProtect adopts a pre-training approach similar to UAP but with a novel mixture-of-perturbation scheme and multi-layer protection loss to enhance protection efficacy. At inference, adaptive targeted protection further boosts protection efficacy with minimal additional cost, and adaptive protection strength improves invisibility.
@@
Universal Adversarial Perturbation (UAP). This concept is introduced in adversarial attacks, through a data-driven approach to find an image-agnostic perturbation from the training dataset [moosavi2017universal]. In our task, we can adapt UAP as in Figure 2b. Given the training dataset $\mathcal{X_{D}}$, we employ a protection loss $\mathcal{L_{P}}$ to derive an universal perturbation $\delta$, constrained within $\eta$-ball to maintain perceptual invisibility.
@@
Upon finalizing the universal perturbation $\delta$ by the training, it can be applied to an image $\mathbf{x}$ to produce a protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$. This is notably practical as it requires no computation at inference. Nevertheless, we found that directly applying UAP to our task significantly compromises protection performance. In Table 1, we compare the protection efficacy of iterative optimization (via PGD) and UAP. We use DISTS [ding2020image] to assess protected images’ quality, while FID [heusel2017gans] evaluates generated image through mimicry by LoRA [hu2021lora] (a higher FID is better protection). As also shown in the qualitative mimicry results by LoRA, UAP leads to a notable degradation in protection efficacy. Hence, it is crucial to refine this pre-training approach, but with a minimal inference cost in a manner akin to UAP.
@@ 3.1 Perturbation Pre-Training
We adopt a simple assignment function $\mathcal{A}$ that can leverage the highly representative latent codes of the VAE encoder. For our training dataset $\mathcal{X}_{D}$, we extract the latent codes $\mathcal{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{D}\}$ using the VAE encoder $\mathcal{E}$. These codes are then clustered into $K$ predefined groups with K-means++ [arthur2007k]. Despite its simplicity, this process shows its robustness empirically, thus we adopt it for our framework.
@@
Training. The assignment function $\mathcal{A}$ is first trained on $\mathcal{X_{D}}$ then MoP is updated on $\mathcal{X_{D}}$ using Eq. 5 by Adam optimizer [kingma2014adam]. All the perturbations ($\delta_{g},\Delta$) are initialized at a resolution of $512\times 512$. When applied to input images of different resolutions, we resize the perturbation through bilinear interpolation. Since two perturbations are used in MoP, $\delta_{g},\Delta$ are constrained within a $(\eta/2)$-ball.
@@ 3.2 Adaptive Inference
Adaptive Targeted Protection. In targeted protection, target image influences protection performance. liang2023mist noted that patterned image is effective, while zheng2023understanding reported that the protection efficacy varies with pattern repetitions. Though they provide valuable insights, they do not consider the relationship between the target and input images. Instead, we investigate such relationship and propose an adaptive targeted protection approach.
@@
In our analysis, we use two target images characterized by either low or high pattern repetition (shown in Suppl. and apply protection to input images with either simple or complex textures. Figure 3 shows mimicry outputs by LoRA of these four protection scenarios. Interestingly, our observations suggest a relationship between the pattern repetition of the target image and the texture complexity of an input image. zheng2023understanding noted that protection efficacy improves as pattern repetition increases but declines when it becomes too high. We suspect that their results might be biased by fixing the input image, thereby keeping the texture factor constant. In contrast, our findings indicate that tailoring the pattern image to match the texture complexity of an input image can improve protection efficacy.
@@
AdvDM [liang2023adversarial] 1210s /   35s 0.197 / 220.0 0.173 / 303.8 0.153 / 357.6 0.271 / 212.5
PhotoGuard [salman2023raising]   370s /    7s 0.203 / 223.0 0.189 / 308.7 0.107 / 350.9 0.209 / 219.1
Anti-DB [van2023anti] 7278s / 225s 0.239 / 214.4 0.162 / 301.4 0.114 / 347.7 0.294 / 225.4
Mist [liang2023mist] 1440s /   40s 0.185 / 217.2 0.154 / 307.5 0.129 / 357.0 0.223 / 223.7
Impasto [ahn2024imperceptible]   830s /   19s 0.201 / 213.8 0.198 / 298.4 0.123 / 352.4 0.207 / 215.5
SDST [xue2023toward] 1410s /   24s 0.242 / 219.2 0.244 / 302.1 0.152 / 354.1 0.237 / 222.7
@@
Adaptive Protection Strength. According to the Weber–Fechner law, humans are more adept at detecting subtle changes in regions with simple textures than in complex ones. Upon this, ahn2024imperceptible adapted human visual system to the image protection task. They utilized multiple perception maps and adjusted perturbation strength during optimization to effectively enhance invisibility. However, the perception maps used by ahn2024imperceptible are computed before the injection of perturbations, leading us to hypothesize that these may not perfectly align with the actual perceived perturbations intended for protection. Moreover, their reliance on a combination of traditional perceptual map algorithms resulted in slow processing and performance limitations. To address these challenges, we first apply MoP to create a surrogate protection image and then generate a perceptual map, which helps resolve the identified issues.
Specifically, we generate a surrogate protection image $\hat{\mathbf{x}}$ through Eq. 6. Subsequently, we create a spatial perceptual map using LPIPS [zhang2018unreasonable], i.e. $\mathbf{M}=\text{LPIPS}(\mathbf{x},\hat{\mathbf{x}})$. As shown in Figure 4, the perceptual map $\mathbf{M}$ is remarkably aligned with human cognitive perspectives. Hence, we utilize this map to produce the final protected image as in below:
@@
Here, $\mathcal{S}(\cdot)$ is a scaling function (more details are in Suppl. and $\mathbf{M}$ serves as a distance map, necessitating an inversion step. Since this process requires only the forward pass of the LPIPS backbone, the additional computational cost at inference is minimal, thus economically enhancing invisibility without significant overhead. In Suppl. we summarize the perturbation pre-training and inference process in the form of an algorithm.

### v2

@@ 1 Introduction
Large-scale diffusion models [50, 16, 43, 41] have achieved remarkable success in the realm of image synthesis task, revolutionizing the way we create and manipulate digital imagery [57, 21, 18, 34, 40]. By adopting recently emerged personalization techniques, it is now possible to develop diffusion models that can generate images in one’s unique styles [44, 11, 24, 49, 2]. However, the very capabilities that make these models so valuable also pose significant risks. Diffusion models can allow malicious users to replicate an individual’s artwork without consent, easily stealing their creative output [48]. Furthermore, diffusion models are exceedingly adept at mimicking faces and objects [44, 24]. This proficiency is not limited to benign applications but can be extended to the making of deepfakes or fake news. The implications for societal harm are profound, as these technologies could be used to undermine public trust, distort political discourse, and violate personal privacy.
@@
Current efforts to protect images against diffusion models have evolved through adversarial perturbation [12]. By introducing perturbations to a given image, they render the protected image resistant to diffusion-based personalization methods, causing the diffusion model to generate distorted outputs. Existing methods primarily rely on two approaches: minimizing the distance between the latent features of the protected and a pre-defined target image within a VAE encoder (e.g. PhotoGuard [46] and Glaze [48]), or fooling the noise prediction network (e.g. AdvDM [28] and Anti-DB [51]). Despite numerous advancements, all current frameworks fundamentally start from an approach that optimizes perturbations during inference [33] when protection is requested, casting doubt on the practicality of these methods to be used in many real applications.
@@
For a protection framework to become a practical solution, the following conditions should be satisfied: (1) Protection Efficacy: Obviously, image protection method should effectively counteract against diffusion-based personalization. Many recent studies have aimed to maximize this aspect by analyzing the internal dynamics of diffusion models [56, 61]. (2) Invisibility: Although some studies strive to be as invisible as possible, adversarial perturbations inevitably leave visible traces on images. However, for commercial services, minimizing these traces is crucial. Moreover, as analyzed by Ahn et al. [1], the task of preventing mimicry tends to leave more noticeable traces, particularly in images with flat textures (e.g. cartoon or illustration). Despite its importance, only a handful of studies highlight this issue [1]. (3) Latency: All existing methods update perturbations during the inference phase through iterative optimization. However, this process is inherently time-consuming; e.g. protecting a $512^{2}$-px image can take up to 5-120 minutes on a CPU and 7-200 seconds even on a high-end A100 GPU. The extensive time required by current protection methods poses a significant barrier that prevents ordinary users from utilizing image protection, thereby still exposing them to the risk of misuse of generative models. Consequently, in order to democratize this technology, it is crucial to develop protection framework that can operate effectively on less powerful devices. Unfortunately, despite extensive progresses in the image protection task, there is scant focus on reducing latency.
@@
In this work, we propose a novel protection framework, FastProtect, which meets all the requirements, with a particular focus on latency. FastProtect diverges from prevalent approaches by adopting a pre-trained perturbation. Although universal adversarial perturbation (UAP) [36] is proposed in adversarial attack, we observed that it significantly reduces protection efficacy for our task. To counteract this, we introduce a mixture-of-perturbations (MoP). Unlike UAP, which relies on a single perturbation, MoP prepares multiple perturbations and selects one of them based on the input image’s latent code. We also propose a multi-layer protection (MLP) loss that utilizes the intermediate features when calculating the protection loss. It intensifies the distortion when the protected images go through fine-tuning by personalization methods, thus enhancing the protection efficacy without any additional cost at inference.
@@
For the inference, we introduce an adaptive targeted protection to enhance protection efficacy with minimal computational overhead. Our observations suggest that in targeted protection, which minimizes the distance between the input image’s latent code and that of a (pre-defined) target, selecting an appropriate target image significantly influences protection performance. Taking this into account, we adaptively determine the best target image based on the input image’s latent code and then apply the perturbation trained for that specific target. To enhance invisibility, we propose an adaptive protection strength approach using the LPIPS [60] distance. Unlike Ahn et al. [1] that leverages multiple just-noticeable difference (JND) maps, our approach does not require ad-hoc modules nor does it significantly increase computational load; it only involves an additional forward pass of LPIPS (i.e. AlexNet [23]).
@@ 2 Background
Diffusion Models. These have gained prominence due to their ability to generate high-quality images. Latent Diffusion Model (LDM) [43] is particularly highligted for its exceptional quality and efficiency. Within LDM, a VAE encoder $\mathcal{E}$, transforms an input image $\mathbf{x}$ into a latent code $\mathbf{z}=\mathcal{E}(\mathbf{x})$. This code is then reconstructed back into the image domain by a decoder $\mathcal{D}$; $\mathbf{x}^{\prime}=\mathcal{D}(\mathbf{z}^{\prime})$. The diffusion process makes a latent code $\mathbf{z}^{\prime}$ by incorporating external factors, $y$, such as textual prompts. The training of LDM is driven by a denoising loss function at each timestep $t$ as:
@@
where a denoising network $\epsilon_{\theta}$ restores the noised latent code $\mathbf{z}_{t}$ based on the timestep $t$ and a conditioning vector $c(y)$. Recent advancements explore the personalization (or few-shot fine-tuning) of LDM with a few reference images through two primary approaches: textual inversion [11], which utilizes the embedding space of CLIP [42] while keeping the denoising network fixed, and model optimization [44], which directly updates the denoising network.
Protection Against Diffusion-based Mimicry. Current protection frameworks apply adversarial perturbations $\delta$ to image $\mathbf{x}$, producing a protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$ using projected gradient descent (PGD) [33]. With a protection loss $\mathcal{L_{P}}$, current methods obtain the protected image of $i$-th optimization step by a signed gradient ascent with step function sgn and step length $\alpha$ as given by
@@
For the objective function, semantic or texture losses are mostly used [27]. The semantic loss, $\mathcal{L_{S}}(\mathbf{x})=\mathcal{L_{SD}}(\mathbf{x})$, is designed to disrupt the denoising process of LDM, misleading it to generate samples that deviate from the original images. AdvDM [28] adopts this loss and Anti-DB [51] is improved to be more robust against DreamBooth [44] by incorporating its update process within the optimization. The texture loss, $\mathcal{L_{T}}(\mathbf{x})=-||\mathcal{E}(\mathbf{x})-\mathcal{E}(\mathbf{y})||^{2}_{2}$, aims to pushing the latent code of $\mathbf{x}$ towards the target image’s latent code. Glaze [48] and PhotoGuard [46] belong to this category. Upon these objective functions, We can now define a universal loss as $\mathcal{L_{P}}=\lambda_{\mathcal{S}}\mathcal{L}_{S}+\lambda_{\mathcal{T}}\mathcal{L}_{T}$, where $\{\lambda_{\mathcal{S}},\lambda_{\mathcal{T}}\}\geq 0$ are balancing factors. Mist [27] utilizes this universal loss function to effectively capture the characteristics of both objectives. Diff-Protect [56], building on Mist, introduces a score distillation trick to achieve more efficient protection. Impasto [1] incorporates perceptual-oriented components designed for imperceptible protection.
@@
Since all the current protection frameworks leverage iterative optimization during inference, they can yield image-specific and high-performing perturbations [46, 27, 48, 28, 1, 51, 56, 61, 55, 52, 30]. Despite this advantage, a significant drawback is the substantial time required to operate protection. For instance, to protect a $512^{2}$-px image, even the fastest model (e.g. PhotoGuard [46]) requires around 7 seconds on an A100 GPU and 350 seconds on a CPU. Consequently, we explore alternatives to this iterative process to enhance efficiency without compromising the protection.
@@
Figure 2: Model overview. (a) Current iterative optimization approaches lack a training phase and perform optimization during inference, resulting in extremely slow protection. (b) UAP [36] introduces pre-training of perturbations, but their image-agnostic nature leads to degraded protection efficacy. (c) Combining the advantages of both paradigms, FastProtect adopts a pre-training approach similar to UAP but with a novel mixture-of-perturbation scheme and multi-layer protection loss to enhance protection efficacy. At inference, adaptive targeted protection further boosts protection efficacy with minimal additional cost, and adaptive protection strength improves invisibility.
@@
Universal Adversarial Perturbation (UAP). This concept is introduced in adversarial attacks, through a data-driven approach to find an image-agnostic perturbation from the training dataset [36]. In our task, we can adapt UAP as in Figure 2b. Given the training dataset $\mathcal{X_{D}}$, we employ a protection loss $\mathcal{L_{P}}$ to derive an universal perturbation $\delta$, constrained within $\eta$-ball to maintain perceptual invisibility.
@@
Upon finalizing the universal perturbation $\delta$ by the training, it can be applied to an image $\mathbf{x}$ to produce a protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$. This is notably practical as it requires no computation at inference. Nevertheless, we found that directly applying UAP to our task significantly compromises protection performance. In Table 1, we compare the protection efficacy of iterative optimization (via PGD) and UAP. We use DISTS [10] to assess protected images’ quality, while FID [15] evaluates generated image through mimicry by LoRA [17] (a higher FID is better protection). As also shown in the qualitative mimicry results by LoRA, UAP leads to a notable degradation in protection efficacy. Hence, it is crucial to refine this pre-training approach, but with a minimal inference cost in a manner akin to UAP.
@@ 3.1 Perturbation Pre-Training
We adopt a simple assignment function $\mathcal{A}$ that can leverage the highly representative latent codes of the VAE encoder. For our training dataset $\mathcal{X}_{D}$, we extract the latent codes $\mathcal{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{D}\}$ using the VAE encoder $\mathcal{E}$. These codes are then clustered into $K$ predefined groups with K-means++ [3]. Despite its simplicity, this process shows its robustness empirically, thus we adopt it for our framework.
@@
Training. The assignment function $\mathcal{A}$ is first trained on $\mathcal{X_{D}}$ then MoP is updated on $\mathcal{X_{D}}$ using Eq. 5 by Adam optimizer [22]. All the perturbations ($\delta_{g},\Delta$) are initialized at a resolution of $512\times 512$. When applied to input images of different resolutions, we resize the perturbation through bilinear interpolation. Since two perturbations are used in MoP, $\delta_{g},\Delta$ are constrained within a $(\eta/2)$-ball.
@@ 3.2 Adaptive Inference
Adaptive Targeted Protection. In targeted protection, target image influences protection performance. Liang and Wu [27] noted that patterned image is effective, while Zheng et al. [61] reported that the protection efficacy varies with pattern repetitions. Though they provide valuable insights, they do not consider the relationship between the target and input images. Instead, we investigate such relationship and propose an adaptive targeted protection approach.
@@
In our analysis, we use two target images characterized by either low or high pattern repetition (shown in Appendix) and apply protection to input images with either simple or complex textures. Figure 3 shows mimicry outputs by LoRA of these four protection scenarios. Interestingly, our observations suggest a relationship between the pattern repetition of the target image and the texture complexity of an input image. Zheng et al. [61] noted that protection efficacy improves as pattern repetition increases but declines when it becomes too high. We suspect that their results might be biased by fixing the input image, thereby keeping the texture factor constant. In contrast, our findings indicate that tailoring the pattern image to match the texture complexity of an input image can improve protection efficacy.
@@
AdvDM [28] 1210s /   35s 0.197 / 220.0 0.173 / 303.8 0.153 / 357.6 0.271 / 212.5
PhotoGuard [46]   370s /    7s 0.203 / 223.0 0.189 / 308.7 0.107 / 350.9 0.209 / 219.1
Anti-DB [51] 7278s / 225s 0.239 / 214.4 0.162 / 301.4 0.114 / 347.7 0.294 / 225.4
Mist [27] 1440s /   40s 0.185 / 217.2 0.154 / 307.5 0.129 / 357.0 0.223 / 223.7
Impasto [1]   830s /   19s 0.201 / 213.8 0.198 / 298.4 0.123 / 352.4 0.207 / 215.5
SDST [56] 1410s /   24s 0.242 / 219.2 0.244 / 302.1 0.152 / 354.1 0.237 / 222.7
@@
Adaptive Protection Strength. According to the Weber–Fechner law, humans are more adept at detecting subtle changes in regions with simple textures than in complex ones. Upon this, Ahn et al. [1] adapted human visual system to the image protection task. They utilized multiple perception maps and adjusted perturbation strength during optimization to effectively enhance invisibility. However, the perception maps used by Ahn et al. [1] are computed before the injection of perturbations, leading us to hypothesize that these may not perfectly align with the actual perceived perturbations intended for protection. Moreover, their reliance on a combination of traditional perceptual map algorithms resulted in slow processing and performance limitations. To address these challenges, we first apply MoP to create a surrogate protection image and then generate a perceptual map, which helps resolve the identified issues.
Specifically, we generate a surrogate protection image $\hat{\mathbf{x}}$ through Eq. 6. Subsequently, we create a spatial perceptual map using LPIPS [58], i.e. $\mathbf{M}=\text{LPIPS}(\mathbf{x},\hat{\mathbf{x}})$. As shown in Figure 4, the perceptual map $\mathbf{M}$ is remarkably aligned with human cognitive perspectives. Hence, we utilize this map to produce the final protected image as in below:
@@
Here, $\mathcal{S}(\cdot)$ is a scaling function (more details are in Appendix) and $\mathbf{M}$ serves as a distance map, necessitating an inversion step. Since this process requires only the forward pass of the LPIPS backbone, the additional computational cost at inference is minimal, thus economically enhancing invisibility without significant overhead. In Appendix, we summarize the perturbation pre-training and inference process in the form of an algorithm.